OpenStudy (anonymous):

how do i find the missing terms in a geometric sequence in this problem: ___, ___, 2, ___,___, 54

OpenStudy (cwrw238):

the 3rd term is 2 and the 6th term = 54 3rd term = ar^2 where a = first term and r = common ration 6th term = ar^5 3rd / 6th term = ar^5 / ar^2 = 2/ 54 from this you can find the value of r and then value of a